How do you solve for t in this equation:
(t/E)+1=t/V

Answer by Alan3354(69443)

(t/E)+1=t/V
------------
t/V - t/E = 1
t*(1/V - 1/E) = 1
t = 1/(1/V - 1/E)
===========
You can:
t = 1/((E - V)/VE)
t = VE/(E - V)